OpenStudy (anonymous):

Find the slope of a line passing through the points (1,3) and (6,-2)

OpenStudy (♪chibiterasu):

Well, we plug these into point-slope form and continue to convert it into slope intercept

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(1\quad ,&3)\quad &(6\quad ,&-2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

\[Slope = \frac{ y _{2}-y _{1} }{ x _{2}-x }\]

OpenStudy (♪chibiterasu):

Yeah that's what it looks like :P

OpenStudy (anonymous):

So -1 would be the correct answer?

jigglypuff314 (jigglypuff314):

-1 would be correct, yes :)

OpenStudy (anonymous):

what about ...Find the slope of a line passing through the points (1,12) and (1,-13).

jigglypuff314 (jigglypuff314):

same idea :) plugin x1=1 , y1=12 , x2=1 , y2=-13

OpenStudy (anonymous):

-12/0?

jigglypuff314 (jigglypuff314):

yes and since you cannot divide by a 0 it would be a vertical line because only a vertical line's slope has a denominator of 0 :)

OpenStudy (anonymous):

What if you have -13/-8, can that be the equivalent of positive 13/8?

jigglypuff314 (jigglypuff314):

yes it would :) hey you're getting pretty good at this! :D

OpenStudy (anonymous):

okay new question...What is the slope of a line that does not intersect the y-axis?

jigglypuff314 (jigglypuff314):

only vertical lines do not intersect the y-axis therefore according to what I had said before, the slope's denominator would be a 0

OpenStudy (anonymous):

So x=1 can be a correct answer?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

What about...do the following ordered pairs represent the same point? (1,3) and (3,1).

jigglypuff314 (jigglypuff314):

if you tried graphing them:  so no, they are not the same.

OpenStudy (anonymous):

heres a new one....write the equation of a line in slope-intercept form with the given slope and point P: Slope = 2; P(7,8)

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b m = slope = 2 so you get y = 2x + b then you can temporarily plug in the given point to find b so when given the point (7, 8) you plug in and get 8 = 2(7) + b solve for b and once you get b plug in what you got for b into y = 2x + b